OpenStudy (anonymous):

Simplify if possible: 4 time the square root of 3 plus 4 times the cube root of 3

OpenStudy (kuoministers):

is this the question? \[\huge4\sqrt{3} + 4\sqrt[3]{3}\]

OpenStudy (kuoministers):

\[\huge48^{\frac{ 1 }{ 2 }} + 4(4^{\frac{ 1 }{ 3 }})\]

OpenStudy (anonymous):

how did you get to this answer?

OpenStudy (kuoministers):

THAT is not a answer it is just simplifing it to indices form.

OpenStudy (kuoministers):

its all up to you to figure out the answer :)